If 3 times a number is decreased by 5, the result is 7. Find the number.
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

If 3 times a number is decreased by 5, the result is 7. Find the number.
------------
Let the number be "x".
EQUATION:
3x-5 = 7
3x = 12
x = 4
=========
Cheers,
Stan H.

Answer by checkley71(8403)

3X-5=7
3X=7+5
3X=12
X=12/3
X=4 ANSWER.
PROOF
3*4-5=7
12-5=7
7=7